Todd and Jessalyn are 25, newly married, and ready to embark on the journey of life.   They both plan to retire 45 years from today. Because their budget seems tight right now, they had been thinking that they would wait at least 10 years and then start investing $2400 per year to prepare for retirement. Jessalyn just told Todd, though, that she had heard that they would actually have more money the day they retire if they put $2400 per year away for the next 10 years - and then simply let that money sit for the next 35 years without any additional payments - than they would have if they waited 10 years to start investing for retirement and then made yearly payments for 35 years (as they originally planned to do).  

Please help Todd and Jessalyn make an informed decision:   

Assume that all payments are made at the end of a year, and that the rate of return on all yearly investments will be 8.4% annually.

A. How much money will Todd and Jessalyn have in 45 years if they put away $200 per MONTH at the end of each month for the next 45 years? (Remember to adjust the 8.4% annual rate to a Rate per month!) (Round this rate per month to 5 places past the decimal)

                                                                   example of rounding: .062134 = .06213 or 6.213%

B. If Todd and Jessalyn wait 25 years (after the kids are raised!) before they put anything away for retirement, how much will they have to put away at the end of each year for 20 years in order to have $1,000,000 saved up on the first day of their retirement 45 years from today?
Answer #1

रूहानि o formulae used : Solution o future value of an = P / (1+r) - 1 Annuity P 2 = Periodic sate no. layment peld period be1 facut 28574.43 .084) 08 - 17. let az (1.084 og both sides Logn z 35 log bose Logo = 35 026350] n z Ant(1.225) x 2 .16.79 --सब बाता sit this Now 35 next $ 35428157 it this amout $ 35428-57 for 25 years, Antout of this amonit = will be alter 35 yeantakeup - det xa (10007) 540 log both sides legn a 540 log (1.007) = 540 (0.0029) * = Antill 566) X E (36.81~- ( put the abovefv = o = P {{1+2) - 1] p [(1-084) - 1 084 -CI 84,000 = p [c1.ose 3 2 -1] - 09 bet x = (1,084) log both sides a z 20 log 70